OpenStudy (anonymous):

Which of the following equations has only one solution? x2 - 8x + 16 = 0 x(x - 1) = 8 x2 = 16

myininaya (myininaya):

the first one has one solution of multiplicity 2

OpenStudy (anonymous):

A,B,or C?

myininaya (myininaya):

i do not see A,B, or C i did say first one though

OpenStudy (anonymous):

A.x2 - 8x + 16 = 0 B.x(x - 1) = 8 C.x2 = 16

jhonyy9 (jhonyy9):

A. has one solution

jhonyy9 (jhonyy9):

because b*(2) - 4ac = 0 and hence x_1,_2 = (8+/_ 0)/2 = 4